Letty is 10 years older than Cory who is half as old as Ben. If the total of their ages is 54, how old is Ben.?

Answer by alicealc(293)

Letty's age = L
Cory's age = C
Ben's age = B

L = 10 + C
C = 1/2 * B
so B = 2C

L + C + B = 54
10 + C + C + 2C = 54
10 + 4C = 54
4C = 54 - 10
4C = 44
C = 44/4 = 11

L = 10 + C = 10 + 11 = 21
B = 2C = 2*11 = 22

so, Ben is 22 years old